A  100 W lamp has steady current of 0.83 A in its filaments. How long is required for 1 mol electrons to pass through the lamp?

1
Expert's answer
2021-12-27T08:02:44-0500

Explanations & Calculations


  • The current is given and let's assume the amount does not change.
  • Then you can calculate how much charge does a mole of elections include.

"\qquad\qquad
\begin{aligned}
\small \text{Charge} &=\small 6.023\times10^{23}\times(-1.6\times10^{-19}C)
\end{aligned}"

  • Once this is calculated, you can proceed with the formula "\small Charge = Current\times time" to calculate the time it would take for that amount to flow.

"\qquad\qquad
\begin{aligned}
\small 6.023\times10^{23}\times|-1.6\times10^{-19}C|&=\small 0.83\,A\times t
\end{aligned}"

  • The t in the equation gives the time, you can give it a try.